# Diffusion Transformer Policy

### v1

@@ Appendix A Additional Training and Model Details
Appendix B Additional Details about Pretraining Data
We choose 15 large datasets from Open X-Embodiment Padalkar et al. (2023) as illustrated in Table 7. We mainly follow Team et al. (2024); Kim et al. (2024) to set the weights. Following Kim et al. (2024), we further resize the image to the size of 224.
Fractal Brohan et al. (2022) DobbE Shafiullah et al. (2023) Droid Khazatsky et al. (2024)
16.15 1.94 13.69
Robo Set Kumar et al. (2024) Viola Zhu et al. (2023b) Kuka Kalashnikov et al. (2018)
2.99 1.30 17.47
BridgeV2 Walke et al. (2023) NYU Franka  Cui et al. (2022) Furniture Bench Heo et al. (2023)
21.86 1.14 6.73
StanfordHydra Belkhale et al. (2023) DLR EDAN  Quere et al. (2020) BerkeleyFanuc  Zhu et al. (2023a)
6.11 0.08 1.07
Jaco Dass et al. (2023) LanguageTable Lynch et al. (2023) toto Zhou et al. (2023)
0.67 6.01 2.78
Table 7: The training dataset mixture
@@ Appendix B Additional Details about Pretraining Data
Appendix C Analysis in Calvin
Appendix D Visualization Analysis
D.1 Zero-shot generalization
Figure 6: Visualization of zero-shot genearalizaton of different models. The first raw is Diffusion Transformer Policy(ours). The second raw of demonstration is OpenVLA Kim et al. (2024), the third raw is Octo Team et al. (2024). Our Model can complete the tasks successfully while both of the others fails.
D.2 Failure cases
We visualize some failure cases and make some further research to explore why model fails on the cases. Figure 7 shows the visualizations of some of these cases. We realize that, models can localize the object and move forward to it in nearly all of the cases, but it is in trouble with predicting a correct pose for the end-effector to grasp the object. Most of the cases fails due to the indiscernible differences on the image or the leak of depth information sometimes. We will explore how to increase the fineness of the prediction for the grasp pose and add the depth information to the model in the future.
Figure 7: Failure examples on Real Robot. Models can not predict a totally correct pose for grasping. Many trajectories fail due to indiscernible differences or the lack of depth information.

### v6

@@ Appendix A Additional Training and Model Details
Appendix B Objects and Tasks Description
The object in our experiments is usually small. The banana is 3cm width and 16cm long. The green block is 2cm width and 4cm long. The kiwifruit is very light with 3cm width and about 4.5cm long. The tiny block is only 0.5cm width and 4cm long. please notice the images (224*224) in Figure LABEL:fig:pickup, Figure LABEL:fig:task_longhorizon, Figure LABEL:fig:complex1 are the input for our model. We demonstrate the pick and place task in Figure LABEL:fig:pickup. In particular, we use the teaching strategy for pick and place tasks to collect data in our experiments. Therefore, we can collect data without remote control or mouse. We further use the remote control to collect more complex data. We demonstrate the complex task and long-horizon tasks as follows,
Pick up the bowl and pour the balls into the box. This task requires to pick up the bowl and pour the balls in the bowl into the box. The bowl is much larger than the block in our experiments. Therefore, we find the model is able to achieve better performance for picking up the bowl. Meanwhile, compared to pick and place tasks, the box of this task is fixed. Thus, it is slightly easier.
Open the box and pick up the green block. In this task, we place the box below the gripper with some variations. The variation of this task is less than the pick up tasks.
Stack the bowls. In this task, we stack the two bowls in the side into the center. For this task, we evaluate it with 20 trials with bowl position variation.
Pick up the banana and move it into the pen container. This task aims to pick up the banana and insert into the small pen container. The banana is aroud 3cm width and 16cm long, while the pen container is 65cm wide. Therefore, it is very difficult to put the banana into the pen container. For this task, we evaluate it with 20 trials with different banana places.
Pick up the banana and move it into the pen container. This task aims to pick up the banana and insert into the small pen container. The banana is aroud 3cm width and 16cm long, while the pen container is 65cm wide. Therefore, it is very difficult to put the banana into the pen container. For this task, we evaluate it with 20 trials with different banana places.
Pick up the mark cup and pouring coffee beans from cup into the bowl. This task aims to pick up the mark cup, and then pouring the coffee beans from the cup into the bowl on top of the box. The mark cup is 200g. Therefore, it is very difficult to pick up it successsfully.
Appendix C Analysis and Discussion
C.1 Comparisons and Analysis with SoTA
In our experiments, we find OpenVLA [28] achieves good pick-up performance compared Octo [69]. However, when the task requires learning rotation operation, Octo is able to achieve better performance, e.g., Open the box. We think it is because Octo predicts continuous actions compared to OpenVLA, and the diffusion policy is not sensitive to action normalizations. Differently, OpenVLA requires to discretize the actions dependent on the action statistics. In our experiments, we calculate the statistics from the 10-shot training samples (all tasks) for OpenVLA, Octo, and DiT policy. It is difficult to obtain suitable statistics for Discretization values. However, diffusion policy is able to predict the continuous actions.
C.2 Zero-shot generalization
Figure 7: Convergence Analysis on OXE dataset [7]. The blue line is DiT Policy, and the orange line is Diffusion action head strategy with the same number of parameters.
C.3 Convergence Analysis
Figure 7 presents the convergence between the baseline method and DiT Policy. We can find that the DiT policy achieves clear faster convergence compared to the diffusion action head strategy. We think this also demonstrates the scalibity of DiT Policy.
C.4 Analysis in Calvin
@@
Appendix D Comparisons on More Diffusion Strategies
In our experiments, we implement a strong baseline based on the core idea of diffusion action head. However, it is slightly different from Octo [69] when we predict action chunks. Octo [69] flattens the action chunks into a single vector with a single embedding. For example, if it requires to predict 8 actions, they predict a $8*7=56$ vector. Unlike the Octo-style diffusion action head, we implement a shared diffusion action head as illustrated in Figure 1 in the main paper, which is more effective. Specially, it is more similar to the Diffussion loss [32] and Diffusion Force [10]. We further devise different diffusion heads, including Unet1D, Transformer on Calvin without pretraining.
Table 10 presents the comparisons about different baseline methods. We can find Diffusion Head with action chunks [69] achieves poor performance on Maniskill2. This indicates that the proposed diffusion action head baseline is more effective compared to the action chunks design [69]. Meanwhile, when we increase the blocks of the action head in our baseline, we do not find clean improvement compared to the three layer MLP. Moreoever, the proposed DiT policy achieves clean better performance on Maniskill2, especially on complex YCB tasks.
[…]
Table 10: Comparision with more diffusion baseline methods on Maniskill2 under success rate. SingleYCB indicates PickSingleYCB, ClutterYCB indicates PickClutterYCB, SingleEGAD indicates PickSingleEGAD. Disc ActionHead indicates Discretized Action Head strategy [6], while Diff ActionHead w/ action chunks is totally similar to Octo strategy [69]. Diffusion Action Head is our baseline. Diffusion Action Head w/ 8 MLP blocks is our baseline with 8 blocks MLP action head.
Method All PickCube StackCube SingleYCB ClutterYCB SingleEGAD
Disc ActionHead 30.2% 41.0% 33.0% 22.0% 1.0% 54.0%
Diffusion Head w/ single token action chunks [69] 36.4% 68.0% 51.0% 26.0% 4.0% 33.0%
Diffusion Action Head 58.6% 86.0% 76.0% 37.0% 24.0% 70.0%
Diffusion Action Head w/ 8 MLP blocks 57.6% 87.0% 73.6% 58.3% 20.8% 53.1%
DiT Policy(ours) 65.8% 79.0% 80.0% 62.0% 36.0% 72.0%
[…]
Methods No. Instructions in a Row (1000 chains)
Unet1D head [12] 76.8% 46.5% 28.8% 18.5% 10.0% 1.80
Transformer head [12] 75.8% 44.8% 26.5% 16.5% 8.0% 1.72
8 layers MLP head 69.8% 42.5% 26.3% 16.8% 11.0% 1.66
3 layers MLP head 75.5% 44.8% 25.0% 15.0% 7.5% 1.68
Single token act chunks 56.5% 18.3% 6.0% 2.8% 0.8% 0.84
Ours 89.5% 63.3% 39.8% 27.3% 18.5% 2.38
Appendix E Additional Details about Pretraining Data
We choose 15 large datasets from Open X-Embodiment [50] as illustrated in Table 12. We mainly follow [69, 28] to set the weights. Following [28], we further resize the image to the size of 224.
Table 12: The training dataset mixture
Fractal [6] DobbE [61] Droid [27]
16.15 1.94 13.69
Robo Set [29] Viola [82] Kuka [24]
2.99 1.30 17.47
BridgeV2 [70] NYU Franka  [13] Furniture Bench [20]
21.86 1.14 6.73
StanfordHydra [1] DLR EDAN  [54] BerkeleyFanuc  [81]
6.11 0.08 1.07
Jaco [14] LanguageTable [41] toto [79]
0.67 6.01 2.78
Figure 8: Visualization of zero-shot genearalizaton of different models. The first raw is Diffusion Transformer Policy(ours). The second raw of demonstration is OpenVLA [28], the third raw is Octo [69]. Our Model can complete the tasks successfully while both of the others fails.
Figure 10: Visualized comparison between Diffusion Transformer Policy and Diffusion Action Head strategy on Real Franka Arm (Pick up the green block). The first raw is DiT Policy, while the second raw is the baseline method with Diffusion Action Head.